17 December 2011 UT: LBC and MODS checkout

Closed for ice on the dome. Winds high, up to 27 m/s.

LBC: Ran up LBC and ran 3 sets of 25 biases.
  • 1st set showed horizontal striping on chips 3 and 4, not on chips 1 and 2. The pattern from image to image changed, but it was the same on both chips 3 and 4. On the end of the 2nd set, the pattern has gone away. And on the 3rd, at 10hr UT, the biases look good. IT #3873 submitted with a jpg of one of the striped biases. _Note added 20111218: At start of night, we saw striping in biases taken soon after turning on LBC, and, as yesterday, this diminished over time. But later, after several hours of not using LBC, we note the horizontal striping again and it is intermittent (see lbcr.20111218.094211.fits for example).

MODS: Ran up MODS and did some closed dome checks.
  • Sieve mask tests done
  • Checked mask table and imaged each of the custom masks (script in /scratch/20111217 on obs2).
    • Vflat 3.0 and 2 seconds exposure time is suitable to see mask and ID. Used r_sdss and g_sdss filters.

-- OlgaKuhn - 17 Dec 2011
Topic revision: r2 - 18 Dec 2011, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback